OpenStudy (anonymous):

solve the equation 3^n-12=1/81

myininaya (myininaya):

go nikita!

OpenStudy (nikita2):

3^n = 12 + 1/81 = (12*81 +1)/81. \[\log_{3}3^{n} = \log_{3}((12*81 +1)/81)\] \[n = \log_{3}(81*12 + 1) - 4 \] but \[\log_{3}(81*12 + 1) \] not integer. (

OpenStudy (anonymous):

my options are 9, 8, -8 and 16

myininaya (myininaya):

did you want n-12 to be the exponent?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok put parathesis around you exponents so we know k?

OpenStudy (anonymous):

okay sorry

OpenStudy (nikita2):

so n-12 = -4

myininaya (myininaya):

its cool

myininaya (myininaya):

gj nakita if you solve n-12=-4 for n you will get your answer

myininaya (myininaya):

n=-4+12=8

OpenStudy (nikita2):

really?)

myininaya (myininaya):

lol